A sample of gas occupies 2.00 L with 5.00 moles present. What would happen to the volume if the number of moles is increased to 10.0 moles? The volume will be liters
1
Expert's answer
2019-04-25T07:58:58-0400

According to Avogadroʾs law

V1/n1 = V2/n2

2.00 L / 5.00 moles = V2 / 10.0 moles

V2 = 4.00 L